OpenStudy (anonymous):

How do you evaluate sin(arccos 4/radical 35)? I thought it would be sin (cos^-1(4/radical 35)) and plug this into my calculator but I'm not positive that is correct. My calculator says 0.7368; is this correct?

OpenStudy (anonymous):

\[\sin \left(\cos ^{-1}\left(\frac{4}{\sqrt{35}}\right)\right)=\sqrt{\frac{19}{35}}=0.736788 \]